### v1

# Near-Optimal Adaptive Transforms for LLM Quantization

Abstract
Quantization to low bitwidth is a standard approach for deploying large language models, however, a few extreme weights and activations stretch the dynamic range and reduce the effective resolution of the quantizer. A common mitigation approach is to apply some fixed orthogonal transforms, such as Hadamard matrices, before quantization, which typically reduces the dynamic range. Yet, these transforms ignore the statistics of the data, and their optimality is currently not understood. In this work, we derive, for the first time, closed-form optimal linear blockwise transforms for joint weight-activation quantization using standard data-free quantizers for common numerical formats. Specifically, we provide derivations of the optimal adaptive (data-aware) transforms for round-to-nearest (RTN), AbsMax-scaled block quantizers for both integer and floating-point formats. The resulting construction, which we call WUSH, combines a Hadamard backbone with a data-dependent component based on second-order moments, yielding a non-orthogonal transform that is provably optimal under mild assumptions and remains structured for efficient implementation. Preliminary experimental results show that our approach consistently improves upon the Hadamard transform for common formats.
1 Introduction
1.1 Overview
Quantization of model weights (NEURIPS2022_c3ba4962; frantar2023optq; MLSYS2024_42a452cb) or activations (pmlr-v202-xiao23c; NEURIPS2024_b5b93943) is now a standard tool for shrinking and accelerating large language models (LLMs) (kurtic2025give), making low-precision inference feasible on a wide range of hardware. A central difficulty, however, is that a few extreme “outlier” weights and activations expand the dynamic range and thereby degrade the effective resolution of low-bit representations.
It is known that one way to mitigate these outliers is to apply linear transforms before quantization (jegou2008hamming; jegou2010product); in the case of LLM quantization, this is often in the form of rotations that spread variance more evenly across channels (cheequip2023; pmlr-v235-tseng24a; NEURIPS2024_b5b93943; liu2025spinquant). For LLMs, Hadamard rotations have been remarkably effective, and blockwise variants aligned with quantization groups have also been shown to be useful in practice. The recent work of egiazarian2025bridginggappromiseperformance shows that the block-wise Hadamard transform offers the best empirical performance for quantization among a large set of existing orthogonal transforms. Yet, these transforms are typically fixed and data-agnostic: in particular, the Hadamard transform does not depend on the statistics of the underlying weights or activations. This raises a natural question: if the Hadamard transform is not data-aware, in what sense can it be considered optimal for quantization?
In this work, we address this question by deriving closed-form optimal linear blockwise transforms for joint weight-activation quantization, which are generally non-orthogonal and adaptive, i.e., data-aware. As opposed to prior methods such as SpinQuant (liu2025spinquant) or FlatQuant (sun2024flatquant), our approach is closed-form and calibration-based, as it does not require training or fine-tuning.
We call our construction WUSH, and we show that it is optimal for floating-point (FP) block quantizers and asymptotically optimal for integer (INT) block quantizers. The name WUSH is a mnemonic for its composition. For a layer with blockwise weight $\bm{W}_{\left(i\right)}$ and calibration activation $\bm{X}_{\left(i\right)}$ in the $i$-th block, the WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}$ is constructed as
$\bm{T}_{\mathrm{hsuw}}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top}$ (1)
where $\bm{W}_{\left(i\right)}^{\prime}$ is the Cholesky decomposition of $\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}=\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}$, the matrices $\bm{U}_{\left(i\right)},\bm{S}_{\left(i\right)},\bm{V}_{\left(i\right)}$ are from the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$, and $\bm{H}$ is a Hadamard matrix. We will define it again in a more exhaustive manner in Eq. (6). WUSH is explicitly adapted to minimize loss in second-order statistics while retaining a structure amenable to efficient implementation and providing a principled, data-aware alternative to heuristic Hadamard-style and blockwise rotations.
1.2 Related Work
Post-training quantization and outliers. Generic post-training quantization schemes, such as RTN (round-to-nearest), OBQ (frantar2022obc), and GPTQ (frantar2023optq), are highly sensitive to heavy-tailed outliers in the weights and activations of LLMs because a small number of extreme values determine the quantization scale. As a result, most quantization levels are spent covering rare extremes instead of the bulk of the distribution, which leads to a significant loss of accuracy.
Non-uniform bitwidths and explicit outlier storage. One common strategy to mitigate outliers is to use non-uniform bitwidths. Methods such as LLM.int8() (NEURIPS2022_c3ba4962), SpQR (dettmers2024spqr), and QUIK (ashkboos-etal-2024-quik) explicitly separate and store outliers in higher precision, while HPTQ (chen2025geometryllmquantizationgptq) uses Huffman encoding to compress the outliers and reduce their storage costs. These approaches can achieve high accuracy at a low effective bitwidth, however, the resulting formats are irregular and require specialized kernels and additional engineering effort for efficient deployment.
Transform-based outlier mitigation. Another line of work applies transforms to the weights and activations before quantization to reduce the impact of outliers. SmoothQuant (pmlr-v202-xiao23c) and AWQ (MLSYS2024_42a452cb) rescale channels to stabilize the dynamic ranges of weights and activations. QuIP (cheequip2023) introduces an incoherence processing step. QuIP# (pmlr-v235-tseng24a), QTIP (NEURIPS2024_6de2e84b), and QuaRot (NEURIPS2024_b5b93943) use Hadamard transforms to spread outlier energy across channel dimensions, while SpinQuant (liu2025spinquant) learns rotations optimized on calibration data. However, their transforms are either heuristic, costly to learn, or primarily designed for weight-only quantization, which limits their accuracy or practicality when applied to fast, per-token activation quantization in large models.
Blockwise transforms and emerging FP formats. The recent FP formats, such as MXFP (ocp_mx_2023) and NVFP (nvidia_blackwell_2024), have motivated blockwise transform schemes. shao2025blockrotationneedmxfp4 indicates that blockwise Hadamard transforms can be more effective under these FP formats than full-layer transforms, but egiazarian2025bridginggappromiseperformance highlights the limited gains achievable with Hadamard alone, linking this to the properties of the underlying weight and activation distributions. In contrast, our work derives a closed-form, data-aware optimal blockwise transform (WUSH) and analyzes how such a transform interacts with both FP and INT block quantizers.
1.3 Methodology
Problem setup. We formulate the transformed weight-activation quantization problem as follows. Define $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ as the weight of a linear layer with $d_{\mathrm{in}}$ input channels and $d_{\mathrm{out}}$ output channels. Define $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$ as the calibration input activation with $d_{\mathrm{in}}$ embedding channels and $d_{\mathrm{batch}}$ tokens. Define $q\left(\cdot\right)$ as a quantizer that maps a matrix of continuous values to quantized values (we will specify $q$ later). Let $\bm{T}_{\mathrm{W}},\bm{T}_{\mathrm{X}}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{in}}}$ be the transforms applied to $\bm{W}$ and $\bm{X}$, respectively. For weight-activation quantization, the output activation $\bm{W}^{\top}\bm{X}$ becomes $q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)$. Our goal is to choose $\bm{T}_{\mathrm{W}}$, $\bm{T}_{\mathrm{X}}$ such that the $L_{2}$ output loss
$\ell=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)-\bm{W}^{\top}\bm{X}\right\|_{\mathrm{F}}^{2}$ (2)
is minimized under the given quantizer $q$. The constant $d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}$ is added to simplify the analysis later.
Block-independent constraint. While the weights can be pre-quantized, the activations must be transformed and quantized dynamically at inference time. To reduce the computational overhead and simplify the problem, we constrain the transforms $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ to be block-diagonal, and the quantizer $q$ to be an RTN (round-to-nearest) quantizer with AbsMax (the maximum absolute value within a group) scales. We assume that the quantization is applied to each sub-column vector of shape $d\times 1$ and that the transform block size aligns with the quantization group size $d$ ($d$ is a factor of $d_{\mathrm{in}}$, and $d$ is a power of $2$). Denote
$\bm{T}_{\mathrm{W}}=\operatorname{diag}\left(\bm{T}_{\mathrm{W}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{W}_{\left(d_{\mathrm{in}}/d\right)}}\right)\quad\text{and}\quad\bm{T}_{\mathrm{X}}=\operatorname{diag}\left(\bm{T}_{\mathrm{X}_{\left(1\right)}},\dots,\bm{T}_{\mathrm{X}_{\left(d_{\mathrm{in}}/d\right)}}\right)$ (3)
with $\bm{T}_{\mathrm{W}_{\left(i\right)}},\bm{T}_{\mathrm{X}_{\left(i\right)}}\in\mathrm{R}^{d\times d}$. We partition $\bm{W}^{\top}=\left[\bm{W}_{\left(1\right)}^{\top},\dots,\bm{W}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{W}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ and $\bm{X}^{\top}=\left[\bm{X}_{\left(1\right)}^{\top},\dots,\bm{X}_{\left(d_{\mathrm{in}}/d\right)}^{\top}\right]$ with $\bm{X}_{\left(i\right)}\in\mathrm{R}^{d\times d_{\mathrm{batch}}}$. Then, we can express the output without quantization as $\bm{W}^{\top}\bm{X}=\sum_{i=1}^{d_{\mathrm{in}}/d}\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}$ and the output with weight-activation quantization as
$q\left(\bm{T}_{\mathrm{W}}\bm{W}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{X}\right)=\sum_{i=1}^{d_{\mathrm{in}}/d}q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right).$ (4)
Finally, define the blockwise output loss as
$\ell_{\left(i\right)}=d_{\mathrm{out}}^{-1}d_{\mathrm{batch}}^{-1}\left\|q\left(\bm{T}_{\mathrm{W}_{\left(i\right)}}\bm{W}_{\left(i\right)}\right)^{\top}q\left(\bm{T}_{\mathrm{X}_{\left(i\right)}}\bm{X}_{\left(i\right)}\right)-\bm{W}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}\right\|_{\mathrm{F}}^{2}.$ (5)
We approximate the layerwise loss as $\ell\approx\sum_{i=1}^{d_{\mathrm{in}}/d}\ell_{\left(i\right)}$ and minimize the blockwise loss $\ell_{\left(i\right)}$ independently.
Optimal block-diagonal transforms. We will prove that there exist closed-form optimal transforms for each block under reasonable assumptions and smooth approximations, which are optimal for FP and asymptotically optimal for INT quantizations. Let $\bm{W}_{\left(i\right)}^{\prime},\bm{X}_{\left(i\right)}^{\prime}\in\mathbb{R}^{d\times d}$ be matrices that satisfy $\bm{W}_{\left(i\right)}^{\prime}\bm{W}_{\left(i\right)}^{\prime\top}=d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $\bm{X}_{\left(i\right)}^{\prime}\bm{X}_{\left(i\right)}^{\prime\top}=d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$, respectively. Without loss of generality (any differences will be canceled in later steps), we let $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ be the lower triangular matrices from the Cholesky decomposition of the second moments $d_{\mathrm{out}}^{-1}\bm{W}_{\left(i\right)}\bm{W}_{\left(i\right)}^{\top}$ and $d_{\mathrm{batch}}^{-1}\bm{X}_{\left(i\right)}\bm{X}_{\left(i\right)}^{\top}$. In the case of $\mathrm{rank}\left(\bm{W}_{\left(i\right)}\right)<d$ or $\mathrm{rank}\left(\bm{X}_{\left(i\right)}\right)<d$, we can dampen the diagonal of the second moment before the Cholesky decomposition. Let the orthogonal matrices $\bm{U}_{\left(i\right)},\bm{V}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and the diagonal matrix $\bm{S}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{W}_{\left(i\right)}^{\prime\top}\bm{X}_{\left(i\right)}^{\prime}=\bm{U}_{\left(i\right)}\bm{S}_{\left(i\right)}\bm{V}_{\left(i\right)}^{\top}$. Let $\bm{H}\in\mathbb{R}^{d\times d}$ be a normalized (orthogonal) Hadamard matrix where each element is $\pm d^{-\frac{1}{2}}$. The optimal $\bm{T}_{\mathrm{W}_{\left(i\right)}}$ and $\bm{T}_{\mathrm{X}_{\left(i\right)}}$ are constructed as
${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{V}_{\left(i\right)}^{\top}\bm{X}_{\left(i\right)}^{\prime\top}\quad\text{and}\quad{\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}=\bm{H}\bm{S}_{\left(i\right)}^{-\frac{1}{2}}\bm{U}_{\left(i\right)}^{\top}\bm{W}_{\left(i\right)}^{\prime\top},$ (6)
respectively. Note that ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}={\bm{T}_{\mathrm{hsuw}}^{-\top}}_{\left(i\right)}$  [^1] The -$\top$ superscript notation means $(\cdot)^{-\top}=((\cdot)^{-1})^{\top}=((\cdot)^{\top})^{-1}$. , which can be easily verified by calculating ${\bm{T}_{\mathrm{hsvx}}}_{\left(i\right)}{\bm{T}_{\mathrm{hsuw}}^{\top}}_{\left(i\right)}=\mathbf{I}$ with $\mathbf{I}$ as the identity matrix.
Remark. The Hadamard matrix $\bm{H}$ is the only data-agnostic ingredient in our optimal formulation in Eq. (6), which explains why it has been empirically shown to be one of the most effective data-agnostic orthogonal transforms.
2 Theoretical Derivation
In this section, we focus on the question of finding optimal transforms that minimize the loss of one block $\ell_{\left(i\right)}$. We omit the subscript $\left(i\right)$ to simplify notation.
2.1 Problem Setup and General Proof Approach
Probabilistic reformulation. We reformulate the problem from a probabilistic perspective. We split the matrices $\bm{W}=\left[\bm{w}_{1},\dots,\bm{w}_{d_{\mathrm{out}}}\right]$ and $\bm{X}=\left[\bm{x}_{1},\dots,\bm{x}_{d_{\mathrm{batch}}}\right]$ into columns $\bm{w}_{k},\bm{x}_{k}\in\mathbb{R}^{d}$ and treat the columns as i.i.d. samples from $d$-dimensional independent distributions $\bm{w}\sim\mathcal{D}_{\mathrm{w}}$ and $\bm{x}\sim\mathcal{D}_{\mathrm{x}}$, respectively. Then the matrices $\bm{W}^{\prime}$ and $\bm{X}^{\prime}$ satisfy $\bm{W}^{\prime}\bm{W}^{\prime\top}=\mathbb{E}_{\bm{w}}\bm{w}\bm{w}^{\top}$ and $\bm{X}^{\prime}\bm{X}^{\prime\top}=\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}$, respectively. We can reinterpret the blockwise loss as $\ell=\mathbb{E}_{\bm{w},\bm{x}}\left(q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)-\bm{w}^{\top}\bm{x}\right)^{2}$.
Unbiased quantization. We assume that $q$ is a stochastic and unbiased quantizer, i.e., for a vector $\bm{\alpha}\in\mathbb{R}^{d}$, the quantization error $\varepsilon\left(\bm{\alpha}\right)=q\left(\bm{\alpha}\right)-\bm{\alpha}$ is a random vector with $\mathbb{E}_{\varepsilon\left(\bm{\alpha}\right)}\varepsilon\left(\bm{\alpha}\right)=\bm{0}$. We further constrain $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$ such that $q\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}q\left(\bm{T}_{\mathrm{X}}\bm{x}\right)$ is unbiased with respect to $\bm{w}^{\top}\bm{x}$. Then, using the first-order approximation, we can split $\ell$ into two non-negative terms:
$\ell=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\left(\bm{T}_{\mathrm{W}}\bm{w}+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right)^{\top}\left(\bm{T}_{\mathrm{X}}\bm{x}+\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)-\bm{w}^{\top}\bm{x}\right)^{2}$ (7)
$=$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)+\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
$\approx$ $\ \mathbb{E}_{\bm{w},\bm{x},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right),\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left(\bm{x}^{\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)+\bm{w}^{\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right)^{2}$
$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{X}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{W}}^{\top}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right)$
$=$ $\ \left(\mathbb{E}_{\bm{w},\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)}\left\|\bm{X}^{\prime\top}\bm{T}_{\mathrm{W}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{W}}\bm{w}\right)\right\|^{2}\right)+\left(\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}\right).$
Problem reduction. We can compute the optimal $\bm{T}_{\mathrm{W}}$ and $\bm{T}_{\mathrm{X}}$ individually for each of the non-negative terms and verify if $\bm{T}_{\mathrm{W}}=\bm{T}_{\mathrm{X}}^{-\top}$. The two terms have similar forms and can be optimized in similar ways. We focus on finding the optimal $\bm{T}_{\mathrm{X}}$ for the second term. Define the $d$-dimensional random variable $\bm{y}=\bm{W}^{\prime\top}\bm{x}\sim\mathcal{D}_{\mathrm{y}}$. Define the transformation matrix $\bm{T}=\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}$. Then, the second term becomes
$\mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\bm{W}^{\prime\top}\bm{T}_{\mathrm{X}}^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{x},\varepsilon\left(\bm{T}_{\mathrm{X}}\bm{x}\right)}\left\|\left(\bm{T}_{\mathrm{X}}\bm{W}^{\prime-\top}\right)^{-1}\varepsilon\left(\bm{T}_{\mathrm{X}}\left(\bm{W}^{\prime-\top}\bm{y}\right)\right)\right\|^{2}$ (8)
$=$ $\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$
Therefore, the two-sided quantization problem is reduced to finding the optimal $\bm{T}\in\mathbb{R}^{d\times d}$ that minimizes the one-sided quantization loss $\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}$ and compute $\bm{T}_{\mathrm{X}}=\bm{T}\bm{W}^{\prime\top}$.
Transformation parameterization. Let the unknown orthogonal matrices $\bm{U}^{\prime},\bm{R}\in\mathbb{R}^{d\times d}$ and the unknown diagonal matrix $\bm{S}^{\prime}\in\mathbb{R}^{d\times d}$ be the singular value decomposition (SVD) of $\bm{T}\bm{U}\bm{S}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}$; thus, $\bm{T}$ can be parameterized as $\bm{T}=\bm{U}^{\prime}\bm{S}^{\prime}\bm{R}^{\top}\bm{S}^{-1}\bm{U}^{\top}$. Denote $\bm{S}=\operatorname{diag}\left(s_{1},\dots,s_{d}\right)$ and $\bm{S}^{\prime}=\operatorname{diag}\left(s^{\prime}_{1},\dots,s^{\prime}_{d}\right)$. Without loss of generality, assume $s_{1}\geq\dots\geq s_{d}>0$ and $s^{\prime}_{1}\geq\dots\geq s^{\prime}_{d}>0$.
Theorem 1 (Optimal Transform)
The optimal configuration of $\bm{T}$ for floating-point (FP) data types is $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$. For integer (INT), the same configuration is optimal up to a $d^{o\left(1\right)}$ factor for zero-mean Gaussian/Laplacian distributed data and within a $d$ factor for any distribution.
Proof We provide detailed proofs for FP and INT in Sections˜2.3 and 2.2, respectively. For each data type, we apply a two-step proof strategy. We first provide a smooth modeling of the quantizer and treat the type casting errors as random noise. Secondly, we calculate the expected quantization error with respect to the parameterized transform and minimize this expectation by solving the corresponding optimization problem.
Some useful notations and basic results are as follows:
  • Denote the $k$-th standard basis vector as $\mathbf{e}_{k}\in\left\{0,1\right\}^{d}$, where the $k$-th element is $1$ and 0 otherwise.
  • The second moment of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}=\bm{W}^{\prime\top}\left(\mathbb{E}_{\bm{x}}\bm{x}\bm{x}^{\top}\right)\bm{W}^{\prime}=\bm{W}^{\prime\top}\bm{X}^{\prime}\bm{X}^{\prime\top}\bm{W}^{\prime}=\bm{U}\bm{S}^{2}\bm{U}^{\top}$.
  • The second moment of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\bm{T}\bm{y}\left(\bm{T}\bm{y}\right)^{\top}=\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}=\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}$.
  • The expected squared norm of $\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}^{2}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$.
  • Similarly, the expected squared norm of $\bm{T}\bm{y}$ is $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$.
  • Because $\bm{S}\succ\bm{0}$, $\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\operatorname{tr}\left(\bm{S}^{2}\right)$. By Jensen’s inequality, $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$.
2.2 Proof for Floating-Point (FP) Types
2.2.1 Quantization Error Modeling
Theorem 2 (FP Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\operatorname{diag}\left(\bm{\eta}\right)\bm{\alpha}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof The FP format $\mathrm{E}a\mathrm{M}b$ is defined as a tuple $\left(s,e,m\right)$ with a sign $s\in\left\{0,1\right\}$, an exponent $e\in\left\{0,\dots,2^{a}-1\right\}$, and a mantissa $m\in\left\{0,\dots,2^{b}-1\right\}$. Assume we do not need to consider the subnormal and special number cases. The tuple $\left(s,e,m\right)$ represents the number $x_{\mathrm{E}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}\left(1+2^{-b}m\right)$.
Using the approximation $1+\left(\cdot\right)\approx 2^{\left(\cdot\right)}$ for $2^{-b}m\in\left[0,1\right]$, we define a smoothed version of $\mathrm{E}a\mathrm{M}b$, named $\mathrm{SE}a\mathrm{M}b$, as $x_{\mathrm{SE}a\mathrm{M}b}=\left(-1\right)^{s}2^{e-2^{a-1}+1}2^{2^{-b}m}=\left(-1\right)^{s}2^{2^{-b}\left(2^{b}e+m\right)-2^{a-1}+1}$ with $2^{b}e+m=2^{b}\left(\log_{2}\left|x_{\mathrm{SE}a\mathrm{M}b}\right|+2^{a-1}-1\right)$ being a $a+b$ bit unsigned integer (concatenating the binary representations of $e$ and $m$). The minimum of $\left|x_{\mathrm{SE}a\mathrm{M}b}\right|$ is $\min\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{-2^{a-1}+1}$ and the maximum is $\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|=2^{2^{-b}\left(2^{a+b}-1\right)-2^{a-1}+1}=2^{2^{a-1}-2^{-b}+1}$.
Denote $\mathrm{SE}a\mathrm{M}b\left(x\right)$ as the quantizer that casts $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ to the nearest number in the $\mathrm{SE}a\mathrm{M}b$ type. Assume $e\in\left(-\infty,2^{a}-1\right]$ so that the smaller values can be represented more precisely. $\mathrm{SE}a\mathrm{M}b\left(x\right)$ can be approximated using integer rounding in the logarithm space and first-order expansion. For $x\neq 0$,
$\mathrm{SE}a\mathrm{M}b\left(x\right)\approx$ $\ \mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\left(\log_{2}\left|x\right|+2^{a-1}-1\right)\right\rceil-2^{a-1}+1}=\mathrm{sgn}\left(x\right)2^{2^{-b}\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil}$ (9)
$\approx$ $\ \mathrm{sgn}\left(x\right)\left(2^{2^{-b}2^{b}\log_{2}\left|x\right|}+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{2^{-b}2^{b}\log_{2}\left|x\right|}2^{-b}\ln 2\right)$
$=$ $\ x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right).$
For $x=0$, $\mathrm{SE}a\mathrm{M}b\left(0\right)\approx\lim_{x\to 0}x\left(1+\left(\left\lfloor 2^{b}\log_{2}\left|x\right|\right\rceil-2^{b}\log_{2}\left|x\right|\right)2^{-b}\ln 2\right)=0$. We model the casting error $\mathrm{SE}a\mathrm{M}b\left(x\right)-x=\left(\ln 2\right)2^{-b}\xi x$ where $x\in\left[-\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|,\max\left|x_{\mathrm{SE}a\mathrm{M}b}\right|\right]$ and the random variable $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
[…]
$q\left(x\right)=$ $\ \mathrm{SE}a\mathrm{M}b\left(xs^{-1}\right)\mathrm{SE}\tilde{a}\mathrm{M}\tilde{b}\left(s\right)=\left(1+\left(\ln 2\right)2^{-b}\xi\right)xs^{-1}\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)s$ (10)
$=$ $\ \left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)x$
with the independent random variables $\xi,\tilde{\xi}\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
Define $\eta=\left(1+\left(\ln 2\right)2^{-b}\xi\right)\left(1+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}\right)-1$. The quantization error is $q\left(x\right)-x=\eta x$.
$\mathbb{E}_{\eta}\eta=\mathbb{E}_{\xi,\tilde{\xi}}\left(\ln 2\right)2^{-b}\xi+\left(\ln 2\right)2^{-\tilde{b}}\tilde{\xi}+\left(\ln 2\right)^{2}2^{-b-\tilde{b}}\xi\tilde{\xi}=0.$ (11)
2.2.2 Loss Minimization
Using Theorem˜2, the minimization objective is
$\mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}$ (12)
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\operatorname{tr}\left(\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right)$
$=$ $\ \mathbb{E}_{\bm{y},\bm{\eta}}\operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\bm{y}^{\top}\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}^{-\top}\right)$
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\operatorname{diag}\left(\bm{\eta}\right)\right)\bm{T}^{-\top}\right)$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$
where $\odot$ represents elementwise multiplication. The lower bound of the trace term in the loss is
$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)$ (13)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\operatorname{diag}\left(\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{1}\right\|^{2},\dots,\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{d}\right\|^{2}\right)\bm{T}^{-\top}\right)$
$=$ $\ \sum_{j=1}^{d}\sum_{k=1}^{d}\left({\mathbf{e}_{j}{T}^{-1}}\mathbf{e}_{k}\right)^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|{T}^{-1}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$
$=$ $\ \sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (rotation-invariance of $L_{2}$ norm)
$\geq$ $\ \sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Cauchy-Schwarz)
$=$ $\ d\left(d^{-1}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}\right)$
$\geq$ $\ d\left(d^{-1}\sum_{k=1}^{d}\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right)^{2}$ (Jensen)
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{U}^{\prime}\bm{S}^{\prime-1}\bm{R}^{\top}\bm{S}\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\right)\right)^{2}$
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$
The equality can be attained by choosing $\bm{U}^{\prime}=\bm{H}$, $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$, and $\bm{R}=\mathbf{I}$ such that
$\ \operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\sum_{k=1}^{d}\left\|\bm{S}\bm{R}\bm{S}^{\prime-1}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}\left\|\bm{R}\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$ (14)
$=$ $\ \sum_{k=1}^{d}\left(\left\|\bm{S}^{\frac{1}{2}}\bm{H}^{\top}\mathbf{e}_{k}\right\|^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}\left(\mathbf{e}_{j}^{\top}\bm{H}^{\top}\mathbf{e}_{k}\right)^{2}\right)^{2}=\sum_{k=1}^{d}\left(\sum_{j=1}^{d}s_{j}d^{-1}\right)^{2}$
$=$ $\ d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}.$
2.2.3 Discussion
Under our smooth modeling in Theorem˜2, for any orthogonal $\bm{T}$ (including the identity $\bm{T}=\mathbf{I}$ and the Hadamard $\bm{T}=\bm{H}$), the minimization objective
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\operatorname{diag}\left(\bm{\eta}\right)\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\bm{y}^{\top}\bm{T}^{\top}\left(\mathbb{E}_{\bm{\eta}}\mathrm{diag}\left(\bm{\eta}\right)^{2}\right)\bm{T}\bm{y}$ (15)
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{S}^{2}\right)\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq\left(\mathbb{E}_{\eta}\eta^{2}\right)d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right),$
so the orthogonal transforms will not be helpful for reducing the quantization error. Also note that the non-trivial choices $\bm{U}^{\prime}=\bm{H}$ and $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ are both essential for reducing the error by at most $d$ times, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario. Either trivial choices of $\bm{U}^{\prime}=\mathbf{I}$ or $\bm{S}^{\prime}=\mathbf{I}$ will lead to the same suboptimal trace term $\operatorname{tr}\left(\bm{T}^{-1}\left(\left(\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\right)\odot\mathbf{I}\right)\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\right)$, which is the same as the case of orthogonal $\bm{T}$.
2.3 Proof for Integer (INT) Data Types
2.3.1 Quantization Error Modeling
We model the quantization error using pseudo noise similar to DiffQ (defossez2022differentiable).
Theorem 3 (INT Quantization Error Modeling)
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we can model the quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=0$.
Proof Define the scalar quantizer of $b$-bit symmetric integer as $q\left(x\right)=\left(\left\lfloor xs^{-1}\right\rfloor+2^{-1}\right)s$, where $x\in\mathbb{R}$, scale $s\in\mathbb{R}_{\neq 0}$, and $xs^{-1}\in\left[-2^{b-1},2^{b-1}\right)$. To simplify the analysis, we model the quantizer as $q\left(x\right)=\left(xs^{-1}+\xi\right)s=x+s\xi$ with random noise $\xi\sim\mathrm{Uniform}\left(-2^{-1},2^{-1}\right)$.
For a vector $\bm{\alpha}\in\mathbb{R}^{d}$, we use the AbsMax scale $s=2^{1-b}\left\|\bm{\alpha}\right\|_{\infty}$. Define $\eta=2^{1-b}\xi$, then we have $s\xi=\eta\left\|\bm{\alpha}\right\|_{\infty}$. The quantization error $\varepsilon\left(\bm{\alpha}\right)=\left\|\bm{\alpha}\right\|_{\infty}\bm{\eta}$ with $\bm{\eta}=\left[\eta_{1},\dots,\eta_{d}\right]^{\top}\in\mathbb{R}^{d}$ being a random vector of i.i.d. samples $\eta\sim\mathcal{D}_{\mathrm{\eta}}$ and $\mathbb{E}_{\eta}\eta=2^{1-b}\mathbb{E}_{\xi}\xi=0$.
2.3.2 Loss Minimization
Using Theorem˜3, the minimization objective is
$\ \mathbb{E}_{\bm{y},\varepsilon\left(\bm{T}\bm{y}\right)}\left\|\bm{T}^{-1}\varepsilon\left(\bm{T}\bm{y}\right)\right\|^{2}=\mathbb{E}_{\bm{y},\bm{\eta}}\left\|\bm{T}^{-1}\left\|\bm{T}\bm{y}\right\|_{\infty}\bm{\eta}\right\|^{2}$ (16)
$=$ $\ \operatorname{tr}\left(\bm{T}^{-1}\left(\mathbb{E}_{\bm{\eta}}\bm{\eta}\bm{\eta}^{\top}\right)\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\left(\mathbb{E}_{\eta}\eta^{2}\right)\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$
$=$ $\ \left(\mathbb{E}_{\eta}\eta^{2}\right)\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}.$
We can obtain a lower bound for the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{T}^{-1}\bm{T}^{-\top}\right)=\operatorname{tr}\left(\bm{U}\bm{S}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\bm{S}\bm{U}^{\top}\right)=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)$. Because $s_{1}^{2}\geq\dots\geq s_{d}^{2}$ and $s_{1}^{\prime-2}\leq\dots\leq s_{d}^{\prime-2}$, using von Neumann’s trace inequality, we obtain:
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=\operatorname{tr}\left(\bm{S}^{2}\bm{R}\bm{S}^{\prime-2}\bm{R}^{\top}\right)\geq\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right).$ (17)
Equality is attained when $\bm{R}=\mathbf{I}$.
Next, we obtain the lower and upper bounds for the term $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}_{\infty}$. Express the expectations as $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=\mathbb{E}_{\bm{y}}\max_{k}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\geq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ (Jensen’s inequality) and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\sum_{k=1}^{d}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\sum_{k=1}^{d}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$. By the relationships among the mean, maximum, and summation values, we have:
$d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\leq d\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},$ (18)
where $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ and $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\mathbf{e}_{k}^{\top}\bm{T}\left(\mathbb{E}_{\bm{y}}\bm{y}\bm{y}^{\top}\right)\bm{T}^{\top}\mathbf{e}_{k}=\mathbf{e}_{k}^{\top}\bm{U}^{\prime}\bm{S}^{\prime 2}\bm{U}^{\prime\top}\mathbf{e}_{k}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|^{2}$.
Lemma 4 (Maximum Inequalities for Tail-Bounded Distributions)
If $X_{1},\dots,X_{d}\in\mathbb{R}$ are zero-mean Gaussian random variables,
$\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}.$ (19)
If $X_{1},\dots,X_{d}\in\mathbb{R}$ are zero-mean Laplacian random variables,
$\mathbb{E}\max_{k}X_{k}^{2}\leq\left(2^{-1}\left(\ln d\right)^{2}+\ln d+1\right)\max_{k}\mathbb{E}X_{k}^{2}.$ (20)
Both inequalities hold without the need for independence.
Proof We provide proofs for Gaussian and Laplacian distributions in Sections˜A.1 and A.2, respectively.
Using Lemma 4, for a tail-bounded $\mathcal{D}_{\mathrm{y}}$ that is a zero-mean multivariate Gaussian or Laplacian, we can further tighten the upper bound by
$\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\min\left\{d^{o\left(1\right)}\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2},\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}\right\}.$ (21)
When $\bm{U}^{\prime}=\bm{H}$, the marginal second moment becomes $\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}=\left\|\bm{S}^{\prime}\bm{U}^{\prime\top}\mathbf{e}_{k}\right\|=\left\|\bm{S}^{\prime}\bm{H}^{\top}\mathbf{e}_{k}\right\|=\sum_{j=1}^{d}s_{j}^{\prime 2}d^{-1}=d^{-1}\operatorname{tr}\left(\bm{S}^{\prime 2}\right)=d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$, which is equalized for all $k$, and $\max_{k}\mathbb{E}_{\bm{y}}\left(\mathbf{e}_{k}^{\top}\bm{T}\bm{y}\right)^{2}$ is minimized to $d^{-1}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}$. Taken together, by setting $\bm{R}=\mathbf{I}$ and $\bm{U}^{\prime}=\bm{H}$, the term $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}$ is bounded by $\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)$ within a gap of $d$ such that
$d^{-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (22)
And for a tail-bounded $\mathcal{D}_{\mathrm{y}}$, the gap is narrowed to $d^{o\left(1\right)}$ such that
$\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)-1}\operatorname{tr}\left(\bm{S^{2}\bm{S}^{\prime-2}}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right).$ (23)
By the Cauchy-Schwarz inequality,
$\operatorname{tr}\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)\operatorname{tr}\left(\bm{S}^{\prime 2}\right)\geq\left(\operatorname{tr}\left(\left(\bm{S}^{2}\bm{S}^{\prime-2}\right)^{\frac{1}{2}}\left(\bm{S}^{\prime 2}\right)^{\frac{1}{2}}\right)\right)^{2}=\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}$ (24)
and the equality is attained when $\bm{S}^{\prime}\propto\bm{S}^{\frac{1}{2}}$. Without loss of generality, we can choose $\bm{S}^{\prime}=\bm{S}^{\frac{1}{2}}$ to minimize both the lower and upper bounds of the quantization loss.
2.3.3 Discussion
Consider the case of $\bm{T}$ being orthogonal. $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}=d$ and $\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|^{2}=\mathbb{E}_{\bm{y}}\left\|\bm{y}\right\|^{2}=\operatorname{tr}\left(\bm{S}^{2}\right)$. For any distribution $\mathcal{D}_{\mathrm{y}}$, the bounds are $\operatorname{tr}\left(\bm{S}^{2}\right)\leq\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}\leq d\operatorname{tr}\left(\bm{S}^{2}\right)$. For the Hadamard $\bm{T}=\bm{H}$ and a tail-bounded $\mathcal{D}_{\mathrm{y}}$, $\left\|\bm{T}^{-1}\right\|_{\mathrm{F}}^{2}\mathbb{E}_{\bm{y}}\left\|\bm{T}\bm{y}\right\|_{\infty}^{2}=d^{o\left(1\right)}\operatorname{tr}\left(\bm{S}^{2}\right)$ with high probability. Because $\operatorname{tr}\left(\bm{S}^{2}\right)\geq d^{-1}\left(\operatorname{tr}\left(\bm{S}\right)\right)^{2}\geq d^{-1}\operatorname{tr}\left(\bm{S}^{2}\right)$, an orthogonal $\bm{T}$ is suboptimal, and the error bounds of an orthogonal $\bm{T}$ can be at most $d$ times larger than those of the optimal $\bm{T}$, with $\operatorname{tr}\left(\bm{S}\right)\approx s_{1}$ being the extreme case scenario.
3 Experiments
3.1 Algorithm
1
Input: weights $\bm{W}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$, activations $\bm{X}\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{batch}}}$, damping ratio $\lambda\in\mathbb{R}$
Output: WUSH transform ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\in\mathbb{R}^{d\times d}$ and pre-quantized weight $\hat{\bm{W}}_{\left(i\right)}\in\mathbb{R}^{d\times d_{\mathrm{out}}}$ for each block $i=1,\dots,d_{\mathrm{in}}/d$
2
[…]
3
[…]
5for $i\leftarrow 1\ \mathrm{to}\ d_{\mathrm{in}}/d\ \mathrm{(in\ parallel)}$ do
6
7 $j_{1},j_{2}\leftarrow\left(i-1\right)d+1,id$
[…]
8
[…]
9
 $\bm{W}^{\prime}\leftarrow\operatorname{Cholesky}\left(\bm{M}_{\mathrm{W}}\right)$ // $\bm{W}^{\prime}\bm{W}^{\prime\top}=\bm{M}_{\mathrm{W}}$
10
 $\bm{U},\bm{\Lambda}\leftarrow\operatorname{Eigendecomposition}\left(\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}\right)$ // $\bm{U}\bm{\Lambda}\bm{U}^{\top}=\bm{W}^{\prime\top}\bm{M}_{\mathrm{X}}\bm{W}^{\prime}$
11
 ${\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\leftarrow\bm{H}\bm{\Lambda}^{-\frac{1}{4}}\bm{U}^{\top}\bm{W}^{\prime\top}$ // WUSH transform
12
 $\hat{\bm{W}}_{\left(i\right)}=q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}^{-\top}\bm{W}[j_{1}:j_{2},1:d_{\mathrm{out}}]\right)$ // pre-quantize weight block
13
14 end for
15
Algorithm 1 Compute WUSH Transforms and Pre-Quantize Weights
Algorithm˜1 summarizes the procedure used to compute the blockwise transforms and pre-quantized weights for each linear layer. In practice, the required second-order information can be obtained either from the calibration activations or from the Hessian matrix; the latter coincides, up to a constant factor, with the Hessian used in GPTQ. As a result, the calibration pipeline closely follows that of GPTQ: layers are processed sequentially, and after processing a layer, the calibration activations are propagated through the quantized layer to provide the inputs for calibrating the next one.
During inference, the forward pass is calculated as
$\sum_{i=1}^{d_{\mathrm{in}}/d}\hat{\bm{W}}_{\left(i\right)}^{\top}q\left({\bm{T}_{\mathrm{hsuw}}}_{\left(i\right)}\bm{X}_{\left(i\right)}\right).$ (25)
with the symbol $\bm{X}$ being overloaded as the new activations instead of the calibration ones.
3.2 Layer Loss
Table 1: Weight-activation RTN quantization losses with identity (I), random rotation (R), Hadamard (H), WUSH without Hadamard (WUS), and WUSH transforms.
Format Transform Q K V O Gate Up Down
MXFP4 I .01105 .01201 .01068 .00435 .00710 .00656 .00547
R .00761 .00773 .00914 .00384 .00556 .00568 .00407
H .00724 .00720 .00860 .00379 .00545 .00561 .00390
WUS .00627 .00722 .00405 .00357 .00576 .00475 .00446
WUSH .00334 .00334 .00330 .00276 .00449 .00439 .00339
NVFP4 I .00423 .00435 .00437 .00234 .00349 .00341 .00241
R .00498 .00501 .00586 .00254 .00363 .00368 .00266
H .00560 .00562 .00670 .00258 .00371 .00379 .00278
WUS .00226 .00236 .00230 .00200 .00304 .00301 .00223
WUSH .00240 .00244 .00228 .00192 .00309 .00302 .00233
INT4 I .17004 .12277 .01315 .00455 .05585 .00983 .01934
R .00579 .00584 .00696 .00294 .00422 .00429 .00310
H .00557 .00555 .00680 .00286 .00409 .00425 .00303
WUS .21268 .14233 .01067 .00454 .05021 .00742 .01308
WUSH .00239 .00243 .00254 .00210 .00343 .00343 .00255
We report the layerwise weight-activation RTN quantization loss ($L_{2}$ loss normalized by the number of elements) for each linear layer in the 18th transformer block of Qwen3-8B, using 32 calibration samples of sequence length 2048 from the FineWeb-Edu (NEURIPS2024_370df50c) dataset. We compare the WUSH transform with the identity (I), random rotation (R) averaged over 10 runs, Hadamard (H), and WUSH without Hadamard (WUS). Table˜1 summarizes the losses for MXFP4, NVFP4, and INT4 formats, where INT4 denotes 4-bit integers with Gaussian MSE clipping and BF16 scales of group size 32. Across formats, WUSH substantially reduces the loss: WUSH consistently yields the smallest loss for MXFP4 and INT4, while WUSH and WUS are almost equally optimal for NVFP4.
These empirical trends are broadly consistent with our theoretical analysis, suggesting that the approximation assumptions (Section˜2.1) and the smooth error models (Sections˜2.2.1 and 2.3.1) capture the dominant sources of quantization loss in practice. At the same time, MXFP4 behaves as a hybrid between ideal FP and INT quantization: the effective mantissa step changes uniformly, imparting INT-like behavior, especially in the subnormal regime and small exponent ranges. This helps explain why Hadamard transforms still provide measurable gains for the MXFP4 format, despite the ideal FP theory (Section˜2.2.3) predicting no benefit for purely orthogonal transforms.
3.3 LLM Benchmarks
We now present preliminary experiments evaluating the accuracy of models using the WUSH transform, relative to the identity (standard RTN) and Hadamard. We conduct our experiment on the LLama-3.2-3B-Instruct, LLama3.1-8B-Instruct, Qwen3-8B, and Qwen3-14B models. We use Platinumbench (vendrow2025large) and the LM Eval Harness (gao2021framework) for evaluation.
The full results for LM Eval Harness tasks are shown in Table˜2. Here, we observe that WUSH achieves the highest average recovery score among all the methods tested, often by large margins. Remarkably, on large models such as Qwen3-14B, it helps narrow the accuracy gap between MXFP4 and NVFP4 to nearly 0.5%.
We plot the results for Platinumbench in Figures˜3, 4, 1 and 2. Again, we observe that WUSH provides significant accuracy improvements, especially in the case of MXFP.
Table 2: Accuracy results on the LM Eval Harness for round-to-nearest (RTN) quantization of weights and activations with identity (I), Hadamard (H), and WUSH transforms.
Model Format Transform MMLU GSM8K HellaSwag WinoGrande Avg. Recovery%
Llama-3.2-3B-Instruct BF16 - 64.43 78.01 73.42 70.09 71.49
NVFP4 I 61.02 71.04 70.90 66.77 67.43 94.33
H 59.91 64.82 69.77 65.59 65.02 90.95
WUSH 61.25 71.78 71.08 67.18 67.82 94.87
MXFP4 I 56.81 60.80 67.30 64.56 62.37 87.25
H 55.99 61.11 68.36 64.09 62.39 87.27
WUSH 58.45 68.60 69.99 66.54 65.89 92.17
Llama-3.1-8B-Instruct BF16 - 72.76 85.06 80.01 77.90 78.93
NVFP4 I 68.64 80.29 77.97 74.43 75.33 95.44
H 66.96 77.41 77.29 74.35 74.00 93.75
WUSH 68.83 78.57 78.22 75.47 75.28 95.37
MXFP4 I 62.21 69.52 73.76 72.61 69.53 88.08
H 63.03 75.44 75.51 71.90 71.47 90.54
WUSH 66.85 75.16 77.28 73.56 73.21 92.75
Qwen3-8B BF16 - 72.98 90.90 75.52 70.56 77.49
NVFP4 I 70.47 88.40 74.44 69.53 75.71 97.70
H 70.19 86.35 73.02 68.11 74.42 96.03
WUSH 70.86 89.33 74.82 69.79 76.20 98.33
MXFP4 I 67.69 84.23 71.24 67.40 72.64 93.74
H 67.53 87.04 71.48 67.64 73.42 94.75
WUSH 70.29 87.95 74.06 69.50 75.45 97.36
Qwen3-14B BF16 - 77.18 91.96 79.84 74.27 80.81
NVFP4 I 75.20 90.22 77.38 73.32 79.03 97.80
H 74.98 92.04 77.76 72.38 79.29 98.12
WUSH 75.43 91.78 78.82 73.18 79.80 98.75
MXFP4 I 72.92 90.22 76.68 71.51 77.83 96.31
H 73.13 87.41 76.29 71.67 77.13 95.44
WUSH 74.56 91.87 78.18 73.15 79.44 98.30
Figure 1: Comparison of different transforms on Qwen3-8B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 2: Comparison of different transforms on Qwen3-14B for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 3: Comparison of different transforms on Llama3.2-3B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
Figure 4: Comparison of different transforms on Llama3.1-8B Instruct for both NVFP4 and MXFP4 RTN quantization on Platinum benchmark tasks. The left table shows accuracy results across the individual Platinum benchmark tasks, while the right plot shows the average accuracy scores together with their standard deviations for each transform.
4 Future Work
A natural next step is to integrate WUSH with advanced post-training quantization schemes such as GPTQ. In this setting, the weights are updated iteratively through error propagation, while WUSH is constructed from the second-order statistics of the updated weights. Designing an algorithm that jointly updates the weights and the associated blockwise transforms is non-trivial and will require adapting both the optimization schedule and the calibration pipeline.
From a systems perspective, efficient kernel support for WUSH is another important direction. Unlike blockwise Hadamard rotations, which reuse a single transform across all blocks, WUSH generally assigns a distinct transform to each block. This per-block specialization complicates kernel reuse and fusion in existing low-level libraries.
Finally, it will be useful to explore structurally simplified variants of WUSH that trade optimality for lower computational overhead. One promising avenue is to approximate the data-aware component with a diagonal matrix, thereby reducing the cost of applying the transforms at inference time.
5 Conclusion
We have studied blockwise weight-activation quantization for LLMs in the presence of heavy-tailed outliers and have shown that, under mild assumptions, this problem admits closed-form linear transforms for both FP and INT block quantizers. Our construction, WUSH, combines a fixed Hadamard backbone with a data-aware component derived from second-order statistics, yielding a non-orthogonal blockwise transform that is provably optimal and remains amenable to efficient implementation. These results clarify the empirical success of Hadamard rotations and indicate that principled, data-aware block transforms can significantly narrow the gap between low-bit and full-precision models.
Acknowledgment
We would like to thank Tijmen Blankevoort for useful discussions and for suggesting the name WUSH, which is a clear improvement over our previous name (HSUW).
Appendix A Maximum Inequalities for Tail-Bounded Distributions
Lemma 5 (Expectation from the Survival Function)
Let $X\geq 0$ be a random variable with the density function $f$ and the survival function $S(t)=\mathbb{P}\left(X\geq t\right)$ with $t\geq 0$. Let $g:\left[0,\infty\right)\to\mathbb{R}$ be integrable and define $G\left(t\right)=\int_{0}^{t}g\left(u\right)\mathrm{d}u$. Assume that $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$. Then $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
Proof Since $X$ has density $f$ and takes values in $0,\infty)$, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t$. The survival function satisfies $S\left(t\right)=\mathbb{P}\left(X\geq t\right)=\int_{t}^{\infty}f(u)\mathrm{d}u$, so $S$ is differentiable almost everywhere with $S^{\prime}\left(t\right)=-f\left(t\right)$. Hence, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}G\left(t\right)f\left(t\right)\mathrm{d}t=-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t$. We now integrate by parts, using $G^{\prime}\left(t\right)=g\left(t\right)$,
$-\int_{0}^{\infty}G\left(t\right)S^{\prime}\left(t\right)\mathrm{d}t=-\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}+\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t.$ (26)
By definition, $G\left(0\right)=0$, and by assumption $\lim_{t\to\infty}G\left(t\right)S\left(t\right)=0$, the boundary term $\left[G\left(t\right)S\left(t\right)\right]_{0}^{\infty}=0$. Therefore, $\mathbb{E}G\left(X\right)=\int_{0}^{\infty}g\left(t\right)S\left(t\right)\mathrm{d}t$.
A.1 Gaussian Distribution
Definition of sub-Gaussian.
A probability distribution of a random variable $X\in\mathbb{R}$ is called sub-Gaussian with variance proxy $\sigma^{2}$ if $\mathbb{E}\exp\left(\left(X-\mathbb{E}X\right)t\right)\leq\exp\left(2^{-1}\sigma^{2}t^{2}\right)$ for all $t\in\mathbb{R}$, and the smallest such $\sigma^{2}$ is called the optimal variance proxy.
Basic properties of sub-Gaussian distributions.
The variance proxy is larger than or equal to the variance: $\sigma^{2}\geq\mathbb{E}\left(X-\mathbb{E}X\right)^{2}$. For a Gaussian distribution, the optimal variance proxy is the same as the variance.
Chernoff bound on the tail: $\mathbb{P}\left(\left|X-\mathbb{E}X\right|\geq t\right)\leq 2\exp\left(-2^{-1}\sigma^{-2}t^{2}\right)$ for all $t\geq 0$.
Maximum inequality.
Let $X_{1},\dots,X_{d}$ be zero-mean Gaussian random variables with variance $\sigma_{1}^{2},\dots,\sigma_{d}^{2}$. Denote $\sigma_{\star}^{2}=\max_{k}\sigma_{k}^{2}=\max_{k}\mathbb{E}X_{k}^{2}$ as the smallest variance proxy for $X_{1},\dots,X_{d}$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (27)
$\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+4d\int_{t_{\star}}^{\infty}t\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+4d\left[-\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t^{2}\right)\right]_{t_{\star}}^{\infty}$
$=$ $\ t_{\star}^{2}+4d\sigma_{\star}^{2}\exp\left(-2^{-1}\sigma_{\star}^{-2}t_{\star}^{2}\right).$
Splitting at $t_{\star}=\sigma_{\star}\sqrt{2\ln\left(2d\right)}$ gives
$\mathbb{E}\max_{k}X_{k}^{2}\leq 2\sigma_{\star}^{2}\ln\left(2d\right)+2\sigma_{\star}^{2}=\left(2\ln\left(2d\right)+2\right)\max_{k}\mathbb{E}X_{k}^{2}$ (28)
without needing the independence of $X_{k}$. Also, considering the trivial bound $\mathbb{E}\max_{k}X_{k}^{2}\leq\mathbb{E}\sum_{k=1}^{d}X_{k}^{2}=\sum_{k=1}^{d}\mathbb{E}X_{k}^{2}\leq d\max_{k}\mathbb{E}X_{k}^{2}$, then $\mathbb{E}\max_{k}X_{k}^{2}\leq\min\left\{d,\left(2\ln\left(2d\right)+2\right)\right\}\max_{k}\mathbb{E}X_{k}^{2}$.
A.2 Laplacian Distribution
For a random variable $X\sim\mathrm{Laplace}\left(\mu,b\right)$ and $t\geq 0$,
$\mathbb{P}\left(\left|X-\mu\right|\geq t\right)=2\int_{t}^{\infty}2^{-1}b^{-1}\exp\left(-b^{-1}x\right)\mathrm{d}x=\exp\left(-b^{-1}t\right).$ (29)
The variance $\mathbb{E}\left(X-\mathbb{E}X\right)^{2}=2b^{2}$.
Let $X_{1},\dots,X_{d}$ be zero-mean Laplacian random variables with diversity $b_{1},\dots,b_{d}$. Denote $b_{\star}=\max_{k}b_{k}=\sqrt{\frac{1}{2}\max_{k}\mathbb{E}X_{k}^{2}}$. For $t\geq 0$, $\mathbb{P}\left(\left|X_{k}\right|\geq t\right)=\exp\left(-b_{k}^{-1}t\right)\leq\exp\left(-b_{\star}^{-1}t\right)$. Using Lemma 5,
$\ \mathbb{E}\max_{k}X_{k}^{2}=2\int_{0}^{\infty}t\ \mathbb{P}\left(\max_{k}\left|X_{k}\right|\geq t\right)\mathrm{d}t=2\int_{0}^{\infty}t\ \mathbb{P}\left(\bigcup_{k=1}^{d}\left\{\left|X_{k}\right|\geq t\right\}\right)\mathrm{d}t$ (30)
$\leq$ $2\int_{0}^{t_{\star}}t1\ \mathrm{d}t+2\int_{t_{\star}}^{\infty}t\left(\sum_{k=1}^{d}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\left(\max_{k}\mathbb{P}\left(\left|X_{k}\right|\geq t\right)\right)\mathrm{d}t$
$\leq$ $2\int_{0}^{t_{\star}}t\ \mathrm{d}t+2d\int_{t_{\star}}^{\infty}t\exp\left(-b_{\star}^{-1}t\right)\mathrm{d}t=\left[t^{2}\right]_{0}^{t_{\star}}+2d\left[-b_{\star}\left(t+b_{\star}\right)\exp\left(-b_{\star}^{-1}t\right)\right]_{t_{\star}}^{\infty}$
$=$ $\ t_{\star}^{2}+2db_{\star}\left(t_{\star}+b_{\star}\right)\exp\left(-b_{\star}^{-1}t_{\star}\right).$
Splitting at $t_{\star}=b\ln d$ gives
$\mathbb{E}\max_{k}X_{k}^{2}\leq b_{\star}^{2}\left(\left(\ln d\right)^{2}+2\ln d+2\right)=\left(2^{-1}\left(\ln d\right)^{2}+\ln d+1\right)\max_{k}\mathbb{E}X_{k}^{2}$ (31)
without needing the independence of $X_{k}$.